Problem 2.7 Annual deposits caming monthly interest Given: You wish to buy a building with monthly payments of $1500 and a $20,000 downpayment. The bank offers a 30-year loan at 5.75% interest compounded monthly. Find: The maximum building price that you can afford.
Answer #1

Amount can be calculated as below

Interest rate compounded monthly.

Tenure of loan =30 years=360 months

Down Payment + Present Worth of all EMIs=20000+1500(P/A,5.75%/12,360)=20000+1500(171.3955)=$277,093.5

Hence the maximum building price that we cah afford is $277,093.5